Hayhayz (hayhayz):

Which of the following is a polynomial with roots 5, 4i, and −4i? f(x) = x3 − 5x2 + 20x − 16 f(x) = x3 − 5x2 + 16x − 80 f(x) = x3 − 20x2 + 5x − 16 f(x) = x3 − 16x2 + 80x − 5

OpenStudy (anonymous):

give me a sec will help u out

Hayhayz (hayhayz):

Thank you

OpenStudy (anonymous):

B.

Hayhayz (hayhayz):

Can you tell me how you got that?

OpenStudy (anonymous):

okay when we do this just bring down the 5. 4*4=16 so this how we got 16 and we just multiply 5*4*4 which equals 80. does that make sense

Hayhayz (hayhayz):

Ohhh yes that does make sense! Could you help me with another?

OpenStudy (anonymous):

for sure

Hayhayz (hayhayz):

Which of the following is a polynomial with roots 4, 6, and −7? f(x) = x3 − 3x2 − 24x + 42 f(x) = x3 − 3x2 − 46x + 168 f(x) = x3 − 24x2 − 42x + 46 f(x) = x3 − 24x2 − 46x + 168 I was gonna do this one based of the answer you gave me from the last one but I guess you have to do this one differently?

OpenStudy (anonymous):

this is a little different u are still multiplying all the number 4,6 &7 to get 168 but the difference is you will multiply 4*6= 24 so D is our I mean your answer lol

Hayhayz (hayhayz):

I was thinking it was D but I just wanted to see what you said haha thanks so much

OpenStudy (anonymous):

anytime hayz, u sound like a nice girl, don't know why anyone would not want to help u.

Hayhayz (hayhayz):

One more please? haha last one I promise!

OpenStudy (anonymous):

its okay bring it on lol

Hayhayz (hayhayz):

Which of the following represents the zeros of f(x) = x3 − 14x2 + 61x − 84? 7, 4, −3 7, 4, 3 7, −4, −3 7, −4, 3

OpenStudy (anonymous):

wait the prone is written in stander form If the polynomial is written in the standard form f(x) = x^3+ax^2+bx+c then the sum of roots should be -a = +3 and the product of roots should be -c = -168. So a = -3 and c = +168. Therefore the answer is B. sorry

Hayhayz (hayhayz):

thank you!